OpenStudy (theopenstudyowl):

I need help quickly... Will give medal to best answer! Let x be a binomial random variable with n = 15 and p = .5. Using the exact binomial calculation and the normal approximation with the continuity correction, find P(x > 6). .3036, .3036 .849, .745 .6964, .6972 .6964, .745 .1527, .3036

OpenStudy (theopenstudyowl):

@ganeshie8

OpenStudy (perl):

the exact binomial, you can use the binomial formula

OpenStudy (perl):

$$ \Large P(x \leq a)=\sum_{i=0}^{a} {n \choose i}p^i (1-p)^{n-i}$$

OpenStudy (perl):

with me so far?

OpenStudy (theopenstudyowl):

yes

OpenStudy (perl):

and use the fact $$ \Large P( x >a ) = 1 - P( x \leq a ) $$

OpenStudy (theopenstudyowl):

ok thnxs, can u help me with som emore plz

OpenStudy (perl):

ok

OpenStudy (perl):

what did you get for this one?

OpenStudy (perl):

I used my calculator: for exact binomial: 1 - binomcdf( 15,.5, 5) = .69638 for the normal approximation : normalcdf( 6.5, 1E99, 15*.5 , sqrt(15*.5*.5)) = .6972